OpenStudy (anonymous):

What is the solution to the equation 5^x=17 ?

OpenStudy (chmvijay):

xlog5 =log17

sam (.sam.):

With \[a^b=c <> \log(a^b)=\log(c)\] Then use \[\log(a^b)=b \log(a)\] Gives \[b \log(a)=\log(c)\] \[b=\frac{\log(c)}{\log(a)}\]

OpenStudy (anonymous):

would it be ... \[\log_{10} 17+\log_{10} 5\] or \[\frac{ \log_{10}17 }{ \log_{10} 5 }\]

OpenStudy (anonymous):

these are the answer choices..

sam (.sam.):

Yes for\[\frac{ \log_{10}17 }{ \log_{10} 5 }\]

sam (.sam.):

\[\log(x)=\log_{10}(x)\]